OpenStudy (anonymous):

Normal and tangent lines at a point.how to solve the tangent,normal line and slope of normal line?someone pls explain the details to me.

OpenStudy (anonymous):

tangent: take derivative, plug in x - value, that gives slope. use point slope formula

OpenStudy (anonymous):

normal, same thing except this time the slope is the negative reciprocal of the tangent slope

OpenStudy (anonymous):

detailed enough? an example might help

OpenStudy (anonymous):

can you solve this pls make it a detailed. y=x^3-2x=1 at (1,2)

OpenStudy (anonymous):

satellite can you solve this pls make it a detailed. y=x^3-2x=1 at (1,2)

OpenStudy (phi):

Do you know how to take the derivative dy/dx of your equation?

OpenStudy (anonymous):

yup its 3x^2-2 right?

OpenStudy (anonymous):

y=x^3-2x+1 at (1,2)

OpenStudy (phi):

Good. That is the slope of your curve.

OpenStudy (anonymous):

yesh. now replace x by 1 to get your slope, use point slope formula to find the line

OpenStudy (anonymous):

@phi not to be argumentative but \[y'=3x^2-2\] is not a slope

OpenStudy (phi):

At each x point, it's different because the curve is curving... it's not a straight line. So at x =1 you can find the slope using the derivative.

OpenStudy (anonymous):

The Formula of the Tangent is \[t(x) = f(a)+f'(a)(x-a)\] The normal is ortogonal to the constant, then it may have the opposite slope. Then it is \[t(x) = f(a)-f'(a)(x-a)\]

OpenStudy (anonymous):

it is the formula for the slope. the slope is a number. you get it by replacing x by whatever number you have for x

OpenStudy (anonymous):

@aoigetsu that is just not right, sorry.

OpenStudy (anonymous):

It isn't? Why?

OpenStudy (anonymous):

normal line is perpendicular.

OpenStudy (anonymous):

at any rate for this problem the derivative is \[3x^2-2\] put x = 1 get \[3-2=1\] so slope is 1

OpenStudy (anonymous):

you have the point (1,2) the slope is 1, so you can find the equation of the tangent line via \[y-2=x-1\] \[y=x+1\]

OpenStudy (anonymous):

Oh yeah Sorry, bad english Try it out: http://graphsketch.com/ The function is x², the derivate is 2x and the point I choose is x=1 Then the functions that will express this relation using the formulas is f(x) = x² tangent=2x-1 normal = -2x+1

OpenStudy (anonymous):

can you pls show how did you get the tangent and normal line?

OpenStudy (anonymous):

now for normal line in this case it is easy because your slope is just 1, so perpendicular line has slope -1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you have the point. it is (1,2)

OpenStudy (anonymous):

you find the slope by replacing x in the derivative by 1

OpenStudy (anonymous):

you get \[3\times 1^2-2=3-2=1\]

OpenStudy (anonymous):

so your slope is 1. then you use the point slope formula \[y-y_1=m(x-x_1)\] with \[m=1, x_1=1, y_1=2\]

OpenStudy (anonymous):

you get \[y-2=1(x-1)\] \[y-2=x-1\] \[y=x+1\]

OpenStudy (anonymous):

this is a lousy example because both the x coordinate and the slope are 1, so you miss some work you would otherwise have to do

OpenStudy (anonymous):

but that is your answer with probably more steps than necessary but i wanted to write them all out in any case

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

am how do you get the tangent?

OpenStudy (anonymous):

you mean the tangent line or the slope of the tangent line?

OpenStudy (anonymous):

the tangent line

OpenStudy (anonymous):

and the normal line?

OpenStudy (anonymous):

ok lets start again with a different problem. maybe will make it clearer. different but easy

OpenStudy (anonymous):

ok pls

OpenStudy (anonymous):

\[f(x)=x^2-4\] at the point (3,5)

OpenStudy (anonymous):

notice that (3,5) is on the graph because \[f(3)=3^2-4=5\] so really i could have just said the 3, and you could have come up with the 5

OpenStudy (anonymous):

now what is \[f'(x)\]?

OpenStudy (anonymous):

am you did not derive it?

OpenStudy (anonymous):

\[f(x)=x^2-4\] what is \[f'(x)\]?

OpenStudy (anonymous):

we need that one. it is pretty easy in this case yes?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

now i want the slope of the line tangent to the graph of \[y=x^2-4\] at the point (3,5) how do i find it?

OpenStudy (anonymous):

i find it by taking \[f'(3)\]

OpenStudy (anonymous):

\[f'(x)=2\times 3=6\] and that is the slope i am looking for

OpenStudy (anonymous):

so now i want an equation for the line tangent to the graph. i have a point, namely (3,5) and i have a slope, it is 6

OpenStudy (anonymous):

so it is straight forward to find the equation for the line. i use the point slope formula and write \[y-5=6(x-3)\]

OpenStudy (anonymous):

some algebra gives me \[y-5=6x-18\] \[y=6x-13\] and that is it, that is the line tangent to the graph at (3,5)

OpenStudy (anonymous):

that is a worked out procedure for finding the tangent line. now on to the "normal" line. we have done most of the work

OpenStudy (anonymous):

the tangent line has slope 6. the normal line is perpendicular, so it has slope \[-\frac{1}{6}\]

OpenStudy (anonymous):

and again we use the point slope formula to find the equation

OpenStudy (anonymous):

am after the tangent line our instructor get T is it the tangent how do you get it?he then solve for the normal line

OpenStudy (anonymous):

i guess its the equation

OpenStudy (anonymous):

the tangent equation

OpenStudy (anonymous):

ok we have the slope of the tangent, it is 6, so we find the slope of the normal by taking the negative reciprocal. it is \[-\frac{1}{6}\]

OpenStudy (anonymous):

the use the point slope formula again \[y-5=-\frac{1}{6}(x-3)\] etc

OpenStudy (anonymous):

you do not use the tangent line to find the normal line. you use the point - slope formula with the negative reciprocal of the slope of the tangent line

OpenStudy (anonymous):

thank u very much for helping me.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you really helped me a lot

OpenStudy (anonymous):

glad to